# Interpretable Multimodal Learning for Tumor Protein-Metal Binding: Progress, Challenges, and Perspectives

### v1

@@ 4.2 Modality combinations
As shown in Table 3, several unimodal methods have been proposed to investigate protein-metal binding using sequence [47, 85, 86, 48, 90, 41, 91, 94] or structural data [87, 88, 93]. However, reliance on a single modality may limit the ability to capture the full complexity of tumor protein-metal interactions. With the rapid advancement of multimodal fusion, researchers are increasingly recognizing the need to integrate multiple data modalities to improve predictive accuracy, improve model intepretability, and support more comprehensive computational analyses. As illustrated in Table 3, a variety of modality combination schemes exist for protein-metal binding models and protein foundation models. These viable modality fusions offer valuable guidance for the enrichment of tumor protein models. The summary of different modality combinations is outlined below.
@@
Sequence and structure each provide valuable but distinct insights into protein-metal interactions, and the limitations associated with unimodal combinations can be addressed by integrating these two modalities. Sequence data alone may overlook the spatial relationships and binding-site conformations that are critical for accurate metal-binding prediction. Conversely, structural data is often constrained by the limited availability of experimentally validated protein structures, especially for novel or mutated tumor proteins. Greener et al. [100] primarily utilized the metalloprotein sequence data and supplemented it with structural insights derived from the “periodic table” of protein structures, based on Taylor’s work. This approach enabled the generation of synthetic sequences corresponding to new protein topologies.
@@
Sequence and text complement each other by allowing the models to access both primary sequence data and biological contextual information. For example, ProteinChat [96] is trained on over 1.5 million carefully curated triplets (protein, text, answer) from the Swiss-Prot dataset, enabling it to capture protein knowledge that mimics expert reasoning.
@@
Sequences, structure, and text can be effectively combined to address annotation gaps in existing protein databases. Currently, approximately 30% of the proteins in UniProt remain unannotated. ProTrek [97] addresses this gap by leveraging this modality combination to predict and annotate the functions of previously uncharacterized proteins. Natural language models further enhance this process by extracting insights from rich textual descriptions while incorporating learning representations of sequence and structural information. For example, ProteinAligner [98] adopts this approach to identify post-translational modifications, capture protein dynamics under different conditions, and map interaction networks directly from the literature.
@@
Sequence, structure, and protein pocket provide a detailed understanding of site-specific information. Trained with AlphaFold-predicted structures [101], ESMBind [84] addresses the challenge of accurately determining the 3D coordinates of metal ions in protein structures, outperforming models such as LMetalSite and M-Ionic, which do not incorporate protein pocket information. ESMBind enables deep learning-based metal-binding prediction at multiple levels, including sequence-level, residue-level, and atomic-level modeling.
@@
Sequence, structure, protein pocket, and text represent a comprehensive integration of previously discussed data modalities. This multimodal combination enables the capture of both low-level molecular features and high-level biological context. The Oneport model [95] exemplifies this framework by generating unified protein representations that support accurate cross-modal retrieval even between modality pairs not explicitly trained together, demonstrating strong generalization and biological relevance. In the context of tumor protein–metal binding, such multimodal combinations are equally applicable. For example, MCCNN effectively organizes and integrates multimodal data for targeted prediction tasks in this domain.
4.3 Multimodal learning
This subsection introduced commonly used multimodal learning methods in medicine and biology and highlighted their potential to guide research on tumor protein-metal interactions. Based on their architectural design, multimodal fusion techniques are mainly classified into early (data-level), intermediate (feature-level), late (decision-level), or hybrid fusion [102, 103]. An additional perspective is the model types used to learn joint feature representations across modalities [104].
@@ 4.3 Multimodal learning
Classical ML models, such as decision trees, support vector machines, and random forests, have been applied to multimodal tasks with concatenated features from different modalities [105]. While these models are interpretable and computationally efficient, they are limited in learning complex relationships or representations from high-dimensional and unstructured data. Deep learning models can effectively capture complex, nonlinear relationships within and across data modalities. Convolutional neural networks (CNNs) can identify local patterns in structured inputs. In the context of tumor protein–metal binding, CNNs can be adapted to learn from protein sequences, 3D structures, and descriptive textual information independently or in combination [106, 107, 108]. Transformer models [109] use self-attention mechanisms to capture relationships between input features across modalities. ProteinAligner [98] utilizes eight transformer layers to process sequence, structural, and textual data. MetaLATTE [85] further introduces a position-sensitive attention mechanism to predict metal specificity, including zinc, lead, and mercury. Graph neural networks are particularly effective for modeling biological data with intrinsic topological structures, such as molecular interactions and protein conformations. For example, graph convolutional networks (GCNs) [110], which apply convolution operations over graph nodes, have been used to predict residue-level binding affinities between yellow fluorescent proteins and rare earth elements. Multi-omics GCN [111] further demonstrates how GCNs can extract complementary features from multi-omics data (mRNA expression, DNA methylation, and microRNA profiles) for disease classification tasks.
@@
Generative models combine feature learning with reconstruction capabilities and are increasingly used for multimodal representation learning and synthetic data generation. Many generative models are built on encoder–decoder architectures, which transform multimodal inputs into a shared latent space and reconstruct them through a decoder. For example, ESBind [84] uses this architecture to process sequence and structural data for predicting protein–metal binding. Multimodal deep autoencoder [112] learns joint representations by fusing similarities from multiple modalities to support downstream tasks like drug–target interaction. A more advanced variant, Variational Autoencoders [113], incorporates both an inference mechanism and a generative process to model the underlying data distribution and produce compact latent representations of sequence and structural data [100]. In addition to representation learning, generative models can address data scarcity by producing synthetic data, such as generating 3D structures from protein sequences [114] and creating descriptive text to supplement incomplete biological annotations.
Although these multimodal learning methods have not yet been widely applied to tumor protein–metal binding, their use in related biomedical fields underscores their potential. With continued adaptation, these methods could enable deeper biological insights and support more accurate predictions in cancer-related metalloprotein research.
5 Interpretability for tumor protein-metal binding models
@@
As shown in Table 3, current models place limited emphasis on interpretability, particularly in the context of tumor protein–metal binding. We first emphasized the universal importance of interpretability and its specific relevance to tumor protein–metal binding. To address gaps in existing methods, we then organized and presented interpretability approaches aimed at enhancing model transparency and facilitating more meaningful biological insights.
@@ 5.1 Importance of interpretability
Interpretability methods can address these issues by demystifying the decision-making process of models, enabling researchers to verify the biological mechanisms captured by the models and establish a scientific basis for predictions [119]. These methods also help identify and reduce bias, ensuring that the models are fair and reliable [120]. Furthermore, by providing clear and understandable explanations of the behavior of the model, interpretability methods improve reproducibility and facilitate the verification of prediction results [122]. An interpretable model can promote new scientific discoveries by offering detailed insights into the biological processes modeled, thereby advancing scientific progress [123, 124, 125, 126].
@@
Enhancing the interpretability of ML models can help scientists gain deeper insights into the interactions between metal ions and tumor proteins, including binding sites, binding modes, and downstream biological effects. Interpretability techniques help highlight key structural features and amino acid residues, offering precise targets for the design and optimization of anti-tumor drugs. By predicting potential binding relationships between novel tumor proteins and metal ions and by evaluating the reliability of these predictions, interpretable models support more informed decision-making. As a result, they can potentially improve the efficiency of drug development while reducing both experimental costs and time.
5.2 Interpretability analyses
To our knowledge, very few studies, either in tumor protein-metal binding or broader protein-metal binding, focus on model interpretability. Instead, most models demonstrate reliability and superiority indirectly, through performance comparisons with previous models or validation against benchmark datasets. While such evaluations reflect predictive capability, they often overlook interpretability, particularly within a biological context. For example, it often remains unclear whether improved performance stems from a deeper understanding of underlying biological mechanisms or simply from better pattern recognition in the data. A model that performs well on general datasets may not maintain its accuracy or biological relevance when applied to tumor-specific scenarios. To address these challenges, we review existing interpretability methods and their potential to enhance biological understanding and domain-specific reliability, as shown in Fig. 1C.
The existing methods are mainly categorized into two groups: inherently interpretable models and post-hoc interpretive methods [115]. Inherently interpretable models are designed with interpretability and transparency in mind from the outset. Their architecture enables direct insight into the decision-making process, making it easier to understand how predictions are derived. Linear models, for example, use a simple combination of features with coefficients that clearly indicate each feature’s impact. Decision trees classify or regress data via hierarchical rules and nodes, making their process intuitive and easy to understand [127]. Generalized additive models (GAMs) combine non-linear functions of multiple features to capture complex relationships while remaining interpretable, allowing for independent interpretation of each feature’s functional form [128]. Rule-based models use a series of “if-then” rules to trigger predictions, simplifying both understanding and validation [129]. Sparse models limit the number of features with non-zero coefficients, thus highlighting key contributions and streamlining interpretation [130]. Modular models decompose the prediction process into distinct, independent components, enabling detailed analysis of each module as well as the overall model behavior [131].
In contrast, post-hoc interpretability methods aim to uncover the inner workings of models through analytical techniques, making them especially practical for complex “black-box” models. They assess models from various perspectives, each tailored to specific situations and offering distinct benefits. Feature importance techniques, such as Local Interpretable Model-agnostic Explanations (LIME) [122], SHapley Additive exPlanations (SHAP) [132], Integrated Gradients [133], Permutation-based Feature Importance Test [134], and Permutation Feature Importance [135], assist users in comprehending the influence of individual features on predictions. For example, LIME approximates the behavior of a complex model by constructing a simple interpretable model within a local region of the model, which can reveal which features the model focuses on when predicting the binding of a certain tumor protein to a metal, such as specific amino acid sequences or domains of the protein. Example-based methods, including SimplEx [136] and Deep k-Nearest Neighbors (DKNN) [137]. SimplEx approximates the model’s prediction for the current input by selecting a set of examples from the training set and combining them in a weighted manner. DKNN, on the other hand, links the prediction of an example with its nearest neighbors in the latent variable space (latent space). Conceptual methods, such as Concept Activation Vectors (CAV) [138] and Concept Activation Regions (CAR) [139], enable evaluation of whether the model utilizes specific concepts. Input-output analysis methods encompass symbolic regression [140] and counterfactual explanations [141]. Symbolic regression methods are employed to directly acquire succinct and comprehensible mathematical expressions from data. These methods are widely utilized to augment the interpretability of neural networks. Counterfactual explanations elucidate which alterations in input features can result in divergent outcomes, thereby assisting users in comprehending how to modify features to attain distinct predictions.
In conclusion, inherently interpretable models ensure transparency from the design stage and complement post-hoc interpretability methods, enhancing the credibility of ML models. In research on the binding of tumor proteins to metals, modular interpretable models, combined with local interpretability analysis or counterfactual explanations of input-output relationships, hold promising prospects.

### v2

@@ 4.2 Modality combinations
As shown in Table 3, several unimodal methods have been proposed for protein-metal binding prediction using sequence [87, 48, 41, 93] or structural data [103, 90, 95]. However, relying on a single modality limits the ability to capture the full complexity of tumor protein-metal interactions. Integrating multiple data modalities, on the other hand, can improve predictive accuracy and model interpretability, thereby supporting more comprehensive computational analyses. Table 3 presents various modality combinations for protein-metal binding prediction and the related protein foundation model that can enhance and enrich tumor protein-metal binding prediction models.
@@
Sequence and structure each provide valuable but distinct insights into protein-metal interactions. Sequence data alone may overlook the spatial relationships and binding-site conformations that are critical for accurate metal-binding prediction. Conversely, structural data is often constrained by the limited availability of experimentally validated protein structures, especially for novel or mutated tumor proteins. Greener et al. primarily utilized the data of the metalloprotein sequence and supplemented it with structural insights derived from the “periodic table” of protein structures, based on Taylor’s work [104]. This approach enabled the generation of synthetic sequences corresponding to new protein topologies.
@@
Sequence and text complement each other by allowing the models to access both primary sequence data and biological contextual information. For example, ProteinChat [98] is trained on over 1.5 million carefully curated triplets (protein, text, answer) from the Swiss-Prot dataset to capture protein knowledge, mimicking expert reasoning.
@@
Sequence, structure, and text can be effectively combined to address annotation gaps in existing protein databases. Currently, approximately 30% of the proteins in UniProt remain unannotated. ProTrek [99] addresses this gap by leveraging sequence, structure, and text data to predict and annotate the functions of previously unidentified proteins. Natural language models can further enhance this task by extracting insights from rich textual descriptions while incorporating sequence and structural information. For example, ProteinAligner [100] adopts this approach to identify post-translational modifications, capture protein dynamics under different conditions, and map interaction networks directly from the literature.
@@
Sequence, structure, and protein pocket provide a detailed understanding of site-specific information. Trained with structures predicted by AlphaFold [105], ESMBind [86] addresses the challenge of accurately determining the 3D coordinates of metal ions in protein structures, outperforming models such as LMetalSite and M-Ionic, which do not incorporate protein pocket information. ESMBind enables deep learning-based metal-binding prediction at multiple levels, including sequence-level, residue-level, and atomic-level modeling.
@@
Sequence, structure, protein pocket, and text represent a comprehensive integration of the previously discussed data modalities. This multimodal combination enables the capture of both low-level molecular features and high-level biological context. The Oneport model [97] exemplifies this approach by generating unified protein representations that support accurate cross-modal retrieval even between modality pairs not explicitly trained together, demonstrating strong generalization and biological relevance. In the context of tumor protein–metal binding, such multimodal combinations are equally applicable. For example, MCCNN effectively integrates multimodal data for targeted prediction tasks in this domain.
4.3 Learning from multimodal data
To learn from multimodal data, we need a fusion scheme and modality representation learning methods. Multimodal fusion schemes can be categorized into early (data-level), intermediate (feature-level), late (decision-level), or hybrid fusion (mix of the previous three schemes) [106, 107]. In the following, we discuss various approaches to learn feature representations from multimodal data [108].
Classical ML models, such as decision trees, support vector machines, and random forests, have been applied to multimodal tasks with concatenated features from different modalities [109]. While these models are interpretable and computationally efficient, they are limited in learning complex relationships or representations from high-dimensional and unstructured data.
@@ 4.3 Learning from multimodal data
Deep learning models can effectively capture complex, nonlinear relationships within and across data modalities. Convolutional neural networks (CNNs) can capture local patterns in structured inputs. In the context of tumor protein–metal binding, CNNs can be adapted to learn protein sequences, 3D structures, and descriptive textual information separately or in combination [110, 111, 112]. Transformer models [113] use self-attention mechanisms to capture relationships between input features across modalities. ProteinAligner [100] utilizes eight transformer layers to process sequence, structural, and textual data. MetaLATTE [87] further introduces a position-sensitive attention mechanism to predict metal specificity, including zinc, lead, and mercury. Graph neural networks are particularly effective for modeling biological data with intrinsic topological structures, such as molecular interactions and protein conformations. For example, graph convolutional networks (GCNs) have been used to predict residue-level binding affinity between yellow fluorescent proteins and rare earth elements [114], and extract complementary features from multi-omics data (mRNA expression, DNA methylation, and microRNA profiles) for disease classification tasks [115]. Deep learning relies heavily on large quantities of high-quality data. However, annotated data on tumor protein-metal binding, especially for mutated samples, are scarce. For instance, despite the availability of tumor-related datasets, MetalPrognosis [11] and MetalTrans [12] still need to leverage cross-species or general protein data for transfer learning.
@@
Generative models perform feature learning with reconstruction capabilities and are increasingly used for multimodal representation learning and synthetic data generation. Many generative models are built on the encoder–decoder architecture, which transforms multimodal input into a shared latent space and reconstructs it through a decoder. For example, ESBind [86] uses such architecture to learn from sequence and structural data for predicting protein–metal binding. Multimodal deep autoencoder [116] learns joint representations by fusing similarities from multiple modalities to support downstream tasks like drug–target interaction. A more advanced variant, Variational Autoencoders [117], incorporates both an inference mechanism and a generative process to model the underlying data distribution and produce compact latent representations of sequence and structural data [104]. In addition to representation learning, generative models can address data scarcity by producing synthetic data, such as generating 3D structures from protein sequences [118] and creating descriptive text to supplement incomplete biological annotations. However, the biological validity of such generated results can be questionable. For example, the predicted metal coordination geometry may not match the actual bond lengths and coordination numbers [86]. Moreover, generation based on known metal-binding protein information may limit their generalization ability for novel metals (such as rare earth elements) or rare tumor mutations [86].
While these models are not yet widely used for tumor protein–metal binding, their application in related domains demonstrates their potential to enhance predictive performance and uncover biological insights in this area.
5 Interpretability in tumor protein-metal binding models
@@
As shown in Table 3, current models place limited emphasis on interpretability, particularly in the context of tumor protein–metal binding. In the following, we first examine the importance of interpretability and its specific relevance to tumor protein–metal binding, and then explore interpretable ML methods for computational biology and summarize their key limitations.
@@ 5.1 Importance of interpretability
Interpretable ML methods can address these issues by demystifying the modal decision-making process, enabling researchers to verify the biological mechanisms captured by the models and establish a scientific basis for predictions [123]. These methods also help identify and reduce bias, ensuring that the models are fair and reliable [124]. Furthermore, by providing clear and understandable explanations of the behavior of the model, interpretable ML methods improve reproducibility and facilitate the verification of prediction results [126]. An interpretable model can promote new scientific discoveries by offering detailed insights into the underlying processes modeled, thereby advancing scientific progress [127, 128, 129, 130].
@@
Enhancing ML model interpretability can help scientists gain deeper insight into the interactions between metal ions and tumor proteins, including binding sites, binding modes, and downstream biological effects. Interpretability techniques help highlight key structural features and amino acid residues, offering precise targets for the design and optimization of anti-tumor drugs. By predicting potential binding relationships between novel tumor proteins and metal ions and by evaluating the reliability of these predictions, interpretable models support more informed decision-making. As a result, they can potentially accelerate drug development while reducing experimental costs.
Most ML-based studies in tumor protein-metal binding or broader protein-metal binding focus primarily on predictive capability, with limited interpretability studies. Consequently, the mechanisms behind performance improvement or differences are often unclear. To better understand these challenges, we next examine existing interpretable ML methods and their potential to enhance biological understanding and domain-specific reliability (Fig. 1C).
5.2 Interpretable ML methods
Following two recent reviews [119, 123], interpretable ML methods for computational biology can be categorized into two approaches: designing inherently interpretable models (i.e., by-design models) and performing post hoc analysis.
In tumor protein-metal binding research, the focus was on certain specific amino acids, such as histidine (His) and cysteine (Cys), because these amino acids contain side-chain groups that can coordinate with metal ions [131, 132]. By conducting site-directed mutagenesis experiments to alter these amino acids, changes can be observed in metal binding capabilities to identify binding sites. Meanwhile, in biological experiments, techniques such as X-ray crystallography and nuclear magnetic resonance (NMR) can be utilized to determine the three-dimensional structures of proteins for analyzing the structural environments of metal binding sites [133, 134]. For instance, some metal binding sites may be located near hydrophobic pockets of proteins or within specific domains.
Using by-design models or performing post hoc feature importance analysis on “black-box” models can facilitate the investigation of whether prediction models can accurately capture the complex binding structures of specific amino acids and metals, as well as their relationships to predicted binding sites.
5.2.1 By-design models
Transparent models, such as linear models and decision trees, have simple structures that make the relationships between inputs and outputs easy to understand. These models are inherently interpretable by design. For example, logistic regression and linear support vector machines (both are linear classifiers) have been applied to classify the chemical stoichiometry of virus-like particles (VLPs) [135]. The learned model weights reveal which parts of the protein sequences contribute more to the binary classification of 60-mer vs 180-mer. The results also show that key amino acids at the ends of $\beta$-sheets (such as valine and isoleucine) may influence VLP assembly more by stabilizing subunit interactions. This application of interpretable linear models can potentially reduce the need for costly lab testing and guide the rational design of VLPs.
Biologically inspired models incorporate known biological concepts into decision-making [41, 95, 136]. For example, GASS-Metal constructs similarity matrices from homologous proteins with the same Enzyme Commission number, rather than relying on general-purpose ones like Blosum62 [136]. This approach reveals how metal-binding functionality is maintained, e.g., through conservative substitutions that preserve key physicochemical properties, even when there are significant sequence differences. This aligns with the hypothesis of “function-driven structure-sequence coevolution” and enhances the trustworthiness of the model.
To capture spatial information and functional interactions in proteins, attention mechanisms have been increasingly used. Attention-based models can learn to identify specific residues or regions that are most relevant to the prediction task. For example, MetalATTE incorporates position-sensitive attention pooling and rotary position embeddings to enhance identification of critical binding sites [87]. Its attention weights highlighted cysteine residues and neighboring amino acids (e.g., histidine, aspartate) as key contributors to heavy metal binding. Furthermore, this interpretability supports biological design. In one case, substituting cysteine residues with alanine at key binding sites led to predicted and experimentally validated loss of function (i.e., non-binding), consistent with the role of cysteine clusters in metal chelation [137].
5.2.2 Post hoc analysis
While by-design models offer built-in interpretability through their structure or biological grounding, post hoc methods aim to explain the predictions of already-trained models, especially when using complex or black-box architectures.
Post hoc interpretability methods can reveal which features or regions a trained model relies on for its predictions. Gradient-based and perturbation-based approaches, such as integrated gradients, Grad-CAM, SHAP, and LIME, are commonly used for this purpose [138, 139, 140, 126]. These methods can help identify influential sequence features or structural positions after the model has been trained.
For example, the hierarchical graph transformer with contrastive learning applies Grad-CAM to project deep model decisions onto the three-dimensional structures of proteins [141]. By visualizing residue-level contributions, it identifies functional sites and overcomes the “homogenized” representations often seen in traditional approaches. In one case, when predicting the function of a protein (e.g., A0A3P7DWR6), Grad-CAM highlighted DNA binding sites that were highly conserved in homologous proteins (e.g., 5H1C), suggesting these residues are evolutionarily constrained and functionally important.
Similarly, Positional SHAP introduces position-sensitive SHAP values to interpret models at a fine-grained level [142]. This approach not only recovers known biological motifs (such as MHC anchor sites) but also uncovers non-trivial interactions between sequence positions, such as those involving electrostatic effects or spatial dependencies.
5.2.3 Interpretability in multimodal integration
Multimodal information integration can improve model performance, but it also raises new questions about the relative contributions and interactions between different modalities. For instance, is it always beneficial to integrate protein sequence and structural information? How can their respective influences be measured? In some tumor proteins with similar sequences but different structures, structure may dominate the identification of metal-binding sites. Conversely, for proteins with similar structures but low sequence conservation, sequence information might be more informative. These hypotheses require systematic testing, and ablation studies provide a principled approach to exploring them.
Ablation experiments reveal potential causal relationships between inputs and outputs by selectively removing specific features or modalities [48, 86]. ESMBind validates the necessity of integrating sequence and structure by comparing its full model with variants using only sequence (ESM-2) or only structure (ESM-IF) [86]. Its findings also suggest that experimental validation of light metal-binding proteins should prioritize structural dynamics, whereas transition metal binding may rely more on sequence conservation.
From a modeling perspective, predictions about metal-binding sites often rely on learned discriminative criteria, such as statistical thresholds for residue co-occurrence or patterns in local structural geometry. Yet, are these criteria consistent with the biochemical principles underlying metal coordination? Integrating biological prior knowledge into model decision-making, either during training or in model architecture, can help align computational predictions with established biochemical mechanisms.
5.3 Limitations of interpretable ML methods
Although recent years have seen progress in interpretable ML methods for protein research, these approaches continue to face fundamental challenges. Addressing these limitations is key to ensuring that interpretable ML becomes a reliable tool in real-world biological and clinical applications.
From the perspective of method evaluation, the absence of “ground truth explanations” poses a significant obstacle to assessing the biological accuracy and relevance of interpretability outcomes [143]. More importantly, the current lack of a standardized evaluation framework makes it difficult to systematically assess and compare interpretable ML methods across diverse tasks and domains [144, 145].
Interpretable methods themselves also have inherent limitations. Feature importance analyses can highlight statistically significant features; however, these often lack clear biological meaning or relevance [146]. Moreover, the presentation formats of interpretability results, such as SHAP values, often diverge from the cognitive expectations of domain experts, impeding clinical translation and practical utility [147]. In addition, instability observed in methods like SHAP and LIME raises concerns about their reliability [123]. Causal inference approaches, such as ablation studies, can evaluate the necessity of specific features but generally fall short of uncovering mechanistic biological insights at the molecular level [148].
Emerging multimodal foundation models present new challenges, particularly in their tendency to favor textual representations while underutilizing essential features from other modalities such as protein structure and physicochemical properties [149, 150]. This bias may risk distorting the interpretability of such models in biological applications.
As model architectures grow more complex, relying on a single interpretability method has become limited in capturing the full decision-making process [151]. Effectively combining multiple interpretability strategies while preserving their individual strengths remains an open challenge. Furthermore, a major concern in current interpretability research is the selective emphasis on well-performing, locally interpretable decisions, with much less effort devoted to explaining where models generalize poorly [123]. This selective transparency hinders robust evaluation and obstructs the translation of ML models from theoretical promise to clinical or experimental practice.